OpenStudy (anonymous):

Solve the percent problem for the unknown base. 38% of what length is 19 cm?

OpenStudy (anonymous):

19 divided by 38 times 100 when you do something times 38 = 19 that something is the one percent value so 38 X = 19 X = 19/38 and then 100 times one percent is the 100 percent value

OpenStudy (anonymous):

Set up a proportion: 38/100=x/19. Cross multiply and solve for x.